OpenStudy (anonymous):

an objects motion is described by the equation d=3sin(8pi t)-2. the displacement, d, is measured in meters. the time t is measured in seconds. a. what is the object position (in meters)at t=0? b. what is the object maximum displacement(in meters) from its t=0 position? c. how much time(in seconds) is required for one oscillation ? d. what is the frequency(measured in Hz) of this oscillation?

OpenStudy (anonymous):

please help me :[

OpenStudy (anonymous):

a). at d=-2 as t=0 sin0=0 b). for sin max or min values are -1<sin t<1 so putting it in d we get d=3-2=1 or d=-3-2=-5 from mean point t=o d=-2 both of these displacements measure as 3 meter. thus answer is 3 meters.